OpenStudy (anonymous):

Solve check 2/3/y = 6 A.1/9 B.1/4 C.9 D.4

OpenStudy (anonymous):

@CoolDude12

OpenStudy (zpupster):

2/3 y = 6 multpky both sides by 3 2y=6*3 and divide both sides by 2 y = (6*3)/2 do the math then check by placing your answer back in the original eqaution

OpenStudy (anonymous):

\[\frac{ 2 }{ 3}y \]

OpenStudy (anonymous):

both sides of what?

OpenStudy (zpupster):

the equal signs or equation

OpenStudy (anonymous):

so multiply \[\frac{ 2 }{ 3 }y\] = 6

OpenStudy (zpupster):

if wee multiply both sides by 3 we get rid of the denominator of three on the left hand side

OpenStudy (zpupster):

so 2y = 6*3

OpenStudy (anonymous):

\[\frac{ 2 }{ 3 } \times \frac{ 6 }{ 3 }\]

OpenStudy (zpupster):

\[(3)\frac{ 2 }{ 3 }y = 6(3)\]

OpenStudy (zpupster):

the threes cancel

OpenStudy (zpupster):

we are left with\[2y = 6(3)\]

OpenStudy (anonymous):

i think 9

OpenStudy (zpupster):

well place nine back in original equation

OpenStudy (zpupster):

\[\frac{ 2 }{ 3}\times9 = 6\]

OpenStudy (zpupster):

there is your check good job!!